OpenStudy (anonymous):

EASY MATH QUESTION, 9TH GRADE!

OpenStudy (anonymous):

?

OpenStudy (anonymous):

care to share it :)

OpenStudy (anonymous):

WAIT A SEC PLEASE!!!

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

THERE!!! HELP PLEASE!!1

OpenStudy (anonymous):

please help me with this, it's for extra credit that I despartaely need. Like, any ideas at all are realty appericated!!

OpenStudy (anonymous):

anyone help? just a little, so I can understand it?

OpenStudy (anonymous):

Can you help me? :)

OpenStudy (debbieg):

Did you read the problem, carefully, 2 or 3 or 7 times, and carefully jot down what you know and what you are trying to find? :) You have 3 heights: 200, 400 and 600 ft. You are given a formula that gives you the wire length (I'll call it w), depending on the height, h. \(\Large w=\sqrt{h^2+(0.55h)^2}\) So, plug in each h to the formula, and get each w. Then it wants to know the minimum wire length needed is you attach THREE wires at EACH height. so once you have the wire length for each height, you multiply each one by 3, and then add them all up, e.g.: \(\Large 3\cdot w_1+3\cdot w_2+3\cdot w_3\)

OpenStudy (anonymous):

I feel stupid, THANK YOU!!! That is..actually pretty easy, thank you. wow. Thank you again, you just make it easy for me and I think I love you okay? Thank you!!!@DebbieG

OpenStudy (debbieg):

lol it's good to be loved. :) happy to help! :)

OpenStudy (anonymous):

So, the answer is 2,188,200 ? Is that right? @DebbieG

OpenStudy (debbieg):

no that's waaaaaayyyy too big. Did you remember to take the square root?

OpenStudy (anonymous):

I was thinking the same thing lol, and square root? at what part?

OpenStudy (anonymous):

OOHHH, YOU MEAN SQUARE ROOT THE WHOLE THING?!! I'M SORRY, I THINK I GOT WHAT YOU MEAN!! ^^

OpenStudy (debbieg):

Well, not the "whole thing", but the formula for the individual wire lengths IS a square root. So you need the sq root for each individual wire length. Just follow the formula... plug & chug!! lol :)

OpenStudy (anonymous):

OHHHH, each of the heights with the e length together then square root them then times by three and added it all together, is that right?

OpenStudy (anonymous):

im sorry for taking your time T^T

OpenStudy (debbieg):

Use this formula: \(\Large w=\sqrt{h^2+(0.55h)^2}\) do it for each of the 3 heights. Enter it carefully on your calculator... and just get the wire length for each height. Add them together. then multiply by 3. (Alternatively, multiply each by 3 and then add those all together... but probably less work to add and then multiply, rather than vice versa :)

OpenStudy (anonymous):

I got the same answer. :(

OpenStudy (anonymous):

the heights after using the formula, 200 = 52,100 400 = 208,400 600 = 468,900 is this right, first?

OpenStudy (debbieg):

what heights are you talking about?? lol.... you're GIVEN height, you need to use that as h in the formula to get the wire length. For h=200: \(\Large w=\sqrt{200^2+(0.55\cdot 200)^2}=\sqrt{52100}\approx228.25\) OK, now I see where you got the 52100 above.... and your others are correct too. But then TAKE THE SQ ROOT. :) THEN add. THEN multiply by 3.

OpenStudy (debbieg):

Your final answer should be under10,000.... not over 2,000,000. :)

OpenStudy (anonymous):

ok, I think I got it, give me a sec to get this done! *^*

OpenStudy (anonymous):

4,108.53 ????!!! IS THAT RIGHT??? :D

OpenStudy (debbieg):

\(\begin{matrix}\Huge \color{gold}{★★★★★★★★★★}\\ \Huge \color{red}{\text{Great job!}} \\ \Huge \color{red}{\text{Gold stars for you!}}\\ \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix} \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\odot \widehat\odot} \\\small\bigtriangleup \\ \Huge \color{red}\smile\end{matrix}\right)}} \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix}\end{matrix}\)

OpenStudy (anonymous):

WOO-HOOO!!! CHAMPION!!!!!WOO-HOO!!! THANK YOU!!!! THANK YOU OMG YOUR AMAZING!!!

OpenStudy (debbieg):

haha, no problem, happy to help. :)

OpenStudy (anonymous):

Can you help me with one more question? a simpler one this time, juts one, pleeaasseee???

OpenStudy (debbieg):

I REALLY have to get going... I'm sorry.... was only sticking around to see you through this problem. :) if you post it, I'm sure someone will help you out!

OpenStudy (anonymous):

Thank you anyways, for the help sweetie I'll posted it, byyeee!!!